OpenStudy (anonymous):

A. –16 B. 16 C. –4 D. 4

OpenStudy (anonymous):

solve plz i will be fan and give medal

OpenStudy (anonymous):

i like your pic E.ali

OpenStudy (anonymous):

-1/8c=2 So we have : x=2/-1/8 and we have : 2/1 * -8/1 = -16 C=-16 Just this :)

OpenStudy (anonymous):

Thank you !

OpenStudy (anonymous):

so c right and thanks

OpenStudy (anonymous):

have one mor can you help

OpenStudy (anonymous):

Sure !

OpenStudy (anonymous):

hang on you said -16 correct

OpenStudy (anonymous):

Yep !

OpenStudy (anonymous):

wait never mind i was stupid for a sec i had to look again

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

Look : Just divided 12 on -3/7 -3/7m=12 and we have : 12/-3/7 12/1*-7/3=-28 m=-28

OpenStudy (anonymous):

y were you wandering if i was a boy just courious

OpenStudy (anonymous):

and thanks

OpenStudy (anonymous):

You re welcome !

OpenStudy (anonymous):

you can e-mail me at camarolover00@hotmail.com or call me at 623-810-8128

OpenStudy (anonymous):

k !